OpenStudy (anonymous):

Find the derivative of f(x) = -3/x at x = -4.

OpenStudy (anonymous):

sorry can't help you on this one, im only 15 haven't done derivatives

OpenStudy (saifoo.khan):

\[\Large f(x) = -3x^{-1}\]Find the derivative now.

OpenStudy (anonymous):

i will bet you have to do this by hand, not by the power rule

OpenStudy (anonymous):

Use the quotient rule to differentiate:\[\bf if \ f(x)=\frac{ g(x) }{ h(x) } \ then \ f'(x)=\frac{ g'(x)h(x)-g(x)h'(x) }{ [h(x)]^2 }\]After obtaining the derivative, just plugin x = -4 and evaluate. You could also do what @saifoo.khan did and use rational exponents to differentiate more easily using Power rule.

OpenStudy (anonymous):

quotient rule?!

OpenStudy (anonymous):

-3/(-4) like this??

OpenStudy (anonymous):

\[\lim_{x\to -4}\frac{-\frac{3}{x}+\frac{3}{4}}{x+4}\]

OpenStudy (anonymous):

from your previous questions i am assuming that you have to do this by the definition

OpenStudy (anonymous):

3/4 3/16 16/3 4/3

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

i made a typo, it is \[\lim_{x\to -4}\frac{-\frac{3}{x}-\frac{3}{4}}{x+4}\] the answer is easy to come by, i think you need the method

OpenStudy (anonymous):

the easy way to do it is to say the derivative of \(-\frac{3}{x}\) is \(\frac{3}{x^2}\) and then plug in \(x=-4\) to get \(\frac{3}{16}\)

OpenStudy (anonymous):

the method is to compute the limit i wrote above \[\lim_{x\to -4}\frac{-\frac{3}{x}-\frac{3}{4}}{x+4}\]

OpenStudy (anonymous):

so what about this one??? Find the derivative of f(x) = -10x2 + 4x at x = 11. -216 -196 -176 -363

OpenStudy (anonymous):

most of it is algebra simplify the expression to get \[\frac{\frac{-3(x+4)}{4x}}{x+4}\] then cancel the \(x+4\) to get \[\frac{-3}{4x}\] then let \(x=-4\)

OpenStudy (anonymous):

\[f(x) = -10x^2 + 4x\] \[f'(x)=-20x+4\] find \[f'(11)\] is the easy way

OpenStudy (anonymous):

so for the first question its 3/16?? and for the last question its 11???

OpenStudy (anonymous):

for the first question it is \(\frac{3}{16}\) for the last question it is \[f'(11)\] where \(f'(x)=-2x+4\)

OpenStudy (anonymous):

oops i meant \(f'(x)=-20x+4\)

OpenStudy (anonymous):

but thats not one of the answers @satellite73

OpenStudy (anonymous):

the answers are -216 -196 -176 -363

OpenStudy (anonymous):

i will let you compute \[f'(11)=-20\times 11+4\]

OpenStudy (anonymous):

oh ok one sec

OpenStudy (anonymous):

so its -216 @satellite73